# Generative Video Semantic Communication via Multimodal Semantic Fusion with Large Model

### v1

@@ Abstract
Despite significant advancements in traditional syntactic communications based on Shannon’s theory, these methods struggle to meet the requirements of 6G immersive communications, especially under challenging transmission conditions. With the development of generative artificial intelligence (GenAI), progress has been made in reconstructing videos using high-level semantic information. In this paper, we propose a scalable generative video semantic communication framework that extracts and transmits semantic information to achieve high-quality video reconstruction. Specifically, at the transmitter, description and other condition signals (e.g., first frame, sketches, etc.) are extracted from the source video, functioning as text and structural semantics, respectively. At the receiver, the diffusion-based GenAI large models are utilized to fuse the semantics of the multiple modalities for reconstructing the video. Simulation results demonstrate that, at an ultra-low channel bandwidth ratio (CBR), our scheme effectively captures semantic information to reconstruct videos aligned with human perception under different signal-to-noise ratios. Notably, the proposed “First Frame+Desc.” scheme consistently achieves CLIP score exceeding 0.92 at CBR = 0.0057 for SNR $>$ 0 dB. This demonstrates its robust performance even under low SNR conditions.
Index Terms:
video semantic communication, visual compression, generative artificial intelligence (GenAI), large model, diffusion model.
@@ I Introduction
On the other hand, advancements in generative artificial intelligence (GenAI) not only enable the creation of realistic, high-quality content but also enhance transmission efficiency through innovations in semantic communication, optimizing bandwidth usage without sacrificing quality [7, 8, 9]. Early GenAI focused on probabilistic graphical models, such as hidden Markov models [10], which generated data with limited quality and diversity. Variational autoencoders (VAEs) marked a breakthrough by producing more complex and realistic data [11, 12], but they faced distribution mismatch issues. To address this, generative adversarial networks (GANs) were proposed [13, 14], optimizing the generator and discriminator through adversarial loss to improve data quality. GANs have been used after DJSCC at the receiver to enhance perception quality [15, 16]. However, GANs are prone to instability, such as mode collapse, resulting in a lack of diversity in the generated samples.
Recent advancements in diffusion models have showcased their impressive ability to generate high-fidelity data conditioned on semantic prompts. For example, textual descriptions like “Panda plays ukulele at home” can now guide the creation of visually coherent short videos with precise semantic alignment [17, 18]. Building on these models’ strengths in semantic understanding and content authenticity, recent studies [19, 20, 21] have demonstrated ultra-low bitrate image transmissions (e.g., $<$0.1 bit per pixel) by transmitting only highly compressed multi-modal semantics, such as text, edge maps, and embeddings. However, existing diffusion-based approaches for image semantic transmission fail to account for temporal frame correlations, making them insufficient for efficient video semantic communication. Extending this paradigm to video transmission remains under-explored: While studies like [22, 23] investigate diffusion-based inter-frame reconstruction for video compression, their scope is limited to error-free channels, leaving challenges in adapting diffusion models to bandwidth-constrained video communication systems.
@@
Therefore, a significant research gap exists in effectively utilizing diffusion models for video reconstruction under practical wireless transmission channels. Delivering video semantics as conditional signals for diffusion models to achieve high semantic scores remains an open challenge. To fill these gaps, we design a generative video semantic communication (GVSC) framework for transmitting video semantics and reconstructing videos, whereby GenAI large models with the diffusion techniques are utilized to generate video aligned with human perception from video semantics of sketch sequences and text description. Our contributions are summarized as follows:
@@
1) To the best of our knowledge, we are the first to propose the novel GVSC framework for wireless video transmission at ultra-low channel bandwidth ratios (CBR). At the transmitter, our semantic extractor captures key visual and textual semantics from the video. Visual data is transmitted using DJSCC, while textual data is encoded with turbo coding and quadrature amplitude modulation (QAM). These semantics encode critical details such as object location, color, size, and actions. At the receiver, pre-trained large GenAI diffusion models integrate both visual and textual semantics for accurate video reconstruction. GVSC’s robustness ensures effective performance even in resource-constrained environments, maintaining high quality at CBRs as low as $10^{-2}$.
@@ II-B2 Joint Source Channel Coding for Visual Modality
$\mathcal{L}(x,x_{0})=k\cdot\text{MSE}(x,x_{0})+(1-k)\cdot\text{LPIPS}(x,x_{0}),\vspace*{-2mm}$ (1)
@@ II-C2 Diffusion Process
After obtaining the semantic embedding, the video generator utilizes these information as conditions to reconstruct the video at the receiver.
During the generation process, the diffusion model $\epsilon_{\theta}$ combines the embeddings of the visual semantics and the video description semantics to perform denoising. Assuming $z_{t}$ is the latent representation at time step $t$, the generation process is as follows:
@@ III-A Transmission Strategy and Model Configuration
[…]
@@
Sketch+Desc.: To enhance the clarity of visual semantics during transmission, we opt for transmitting only the initial frame’s sketch, thereby concentrating resources on its quality. An initial sketch of the video is transmitted with video description. The sketch is encoded using the DJSCC optimized for sketch transmission. Stable Diffusion 3.5 [30] is used to reconstruct the first frame of the video, after which Open-Sora [31] reconstruct the whole video. We configure the DJSCC-sketch and train the model over an SNR range from 0 to 10 dB. Testing is performed across all the SNRs within this range. The total number of symbols required for this strategy is given by $K=1.5N_{d}+N_{s}$ where $N_{s}$ represents the number of transmitted symbols for the sketch, and $1147.561.5N_{d}=$ 1147.56 symbols are required for the video description. The output size of the DJSCC encoder is $32\times 32\times 2$. Therefore, the resource blocks allocated for the sketch data are calculated as $\frac{32\times 32\times 2}{2}=1024$ i.e., resource blocks for sketch. Thus, the total number of resource blocks required for the “Sketch+Desc.” scheme is:$K=1147.56+1024=2171.56$. Consequently, the CBR for this scheme is $0.001$.
First Frame+Desc.: For acquiring more detailed semantic information, we choose to transmit the first frame with video descriptions, which are transmitted using DJSCC. Open-Sora then utilizes these rich semantic inputs to reconstruct the video. This scheme combines a high-quality visual anchor with textual semantics to detail dynamic content, ensuring the high-quality video reconstruction. For this scheme, we have specifically trained the DJSCC to optimize the transmission of the first frame. The video description is transmitted as “Desc. Only” scheme. This scheme achieve a average CBR of $0.0057$.
@@ III-B Evaluation Metrics
The GVSC Framework adopts the following metrics to evaluate semantic similarity. CLIP score: The CLIP score [32], widely adopted due to its training on large image-text datasets [33, 34, 35, 36], effectively captures high-level semantic features and is suitable for assessing semantic similarity. We use the average CLIP score to compare each generated frame with its corresponding original frame. It effectively measures the semantic similarity of each video frame. The CLIP score focuses on frame-level semantic similarity, ensuring each generated frame matches its original counterpart. BERT score: We design a video semantic consistency check scheme. To assess overall semantic similarity, we use Video-LLaVA to generate captions for both videos and evaluate their similarities with a pretrained BERT model [37]. Given Video-LLaVA’s generative diversity, we sample captions 3 times for each video and average the scores for the overall semantic score. The BERT score evaluates the video-level semantic consistency by comparing generated captions, measuring how well the generated video captures the broader narrative and context. PSNR and SSIM [38]: The peak signal-to-noise ratio (PSNR) and structure similarity index measure (SSIM) are traditional video quality metrics, providing insights into pixel and structure-level fidelity. They measure the peak SNR and structural similarity, respectively.
@@ IV-A1 Dataset
We use the WebVid dataset [39] for both training and testing our model. WebVid is a video dataset crawled from the Internet, including videos and corresponding captions. It encompasses a variety of video scenes, including character movements and natural scenery. We use the captions in the WebVid as description semantic information. Each video has $F$ = 8 frames, and we set the video resolution to 256 $\times$ 256, so the height $H$ and width $W$ of each frame is $H$ = $W$ = 256. For the transmission model, we employ PiDiNet [40] to extract over 11,000 sketches from WebVid for training and 129 videos for testing.
@@ IV-A2 Simulation Parameters
For transmitting sketches, we train the DJSCC for a total of 200 epochs using the Adam optimizer and batchsize is 16. The learning rate is $1\times 10^{-4}$. Following the human assessment simulations, we set $k=0.3$ in (1). We train our transmission model over a SNR range of 0 to 10 dB and test all models within this range. Due to the excessive number of parameters, we freeze the GenAI model. We set the frame rate to 4, the inference time-step to 50, and the random seed to $7$.
@@ IV-B Comparison Schemes
H.264/H.265+LDPC: As comparative schemes, we use H.264/H.265 for source coding, combined with low-density parity-check (LDPC) codes for channel coding and 4-QAM. The quantization parameter of H.264/H.265 is adapted to the same CBR as used in the proposed schemes. We set $\text{CBR}_{\text{H.264}}=0.008$ and $\text{CBR}_{\text{H.265}}=0.008$ in our simulations.
DJSCC-RGB: We use DJSCC with three layers of upsampling and downsampling modules. We maintain the original channel setting (3 channels) to directly transmit the RGB frames of the videos.
@@ IV-C Simulation Results
Visual comparisons of different transmission schemes at SNR = 10 dB and SNR = 0 dB are shown in Fig. 2. The first column on the left displays the original video frames (the 1st and the 8th frames). Subsequent columns present visual outcomes of various methods at SNR = 10 dB. The rightmost part highlights the DVST and “First Frame+Desc.” schemes at SNR = 0 dB, both maintaining a CBR of $10^{-2}$. While H.264 and H.265 still preserve detail, they exhibit noticeable compression artifacts. DVST shows the effective transmission at SNR = 10 dB with high visual quality, while DJSCC-RGB undergoes significant visual degradation due to ultra-low CBR, showing highly distorted frames. The “Desc. Only” scheme preserves basic semantic information but loses structural details, leading to relatively poor video reconstruction. Conversely, the “Sketch+Desc.” and “Sketches+Desc.” schemes excel in retaining both structural and textual semantics, as reflected in descriptions such as “A boy is eating corn”. In GVSC, the “First Frame+Desc.” scheme allocates more resources to the first frame, ensuring semantic consistency and excelling in visual clarity even at SNR = 0 dB, as highlighted in the red circles. This method leverages textual descriptions to guide the dynamic aspects of the scene, providing an edge in realism and robustness in the low SNR range.
Fig. 3 shows CLIP score of different schemes for various SNRs. Traditional codecs, e.g., H.264 and H.265, exhibit a cliff effect for SNR $<$ 5 dB, indicating difficulties in preserving textual semantics under strong noisy conditions. DJSCC-RGB shows gradual improvement with increasing SNR but performs modestly due to noise at ultra-low CBR. The “Desc. Only” scheme, relying solely on text, achieves lower CLIP score due to the lack of structural information but slightly outperforms DJSCC-RGB due to higher resolution. The “Sketch+Desc.” and “Sketches+Desc.” schemes, which integrate both structural and textual semantics, demonstrate superior performance as SNR increases. This highlights the effectiveness of combining sketches with textual descriptions. Due to blurriness, DVST exhibits lower CLIP score at low SNRs but it recovers rapidly with increasing SNR. Notably, the “First Frame+Desc.” scheme consistently outperforms other schemes under all SNR conditions and maintains high CLIP score even at low SNRs. This verifies the robustness of our GVSC framework in ensuring semantic integrity under different SNRs.
@@
Figure 7: CLIP score for various SNRs with different k in “Sketch+Desc.” scheme.
@@
Fig. 4 presents BERT score performance of different schemes for various SNR conditions. Similar to CLIP score, H.264 and H.265 have difficulty in transmitting video data effectively at SNR $<$ 5 dB. The BERT scores for “Sketch+Desc.” and “Sketches+Desc.” schemes are lower than those of H.264 and H.265 because traditional schemes transmit detailed visual information closely related to the original video, leading to higher text similarity scores. When generating descriptions, the visual model might include non-critical details such as hair color or eye direction, which are not always relevant in the video generation context. Since BERT focuses on specific textual details, even slight discrepancies can result in lower scores. From Fig. 4, we can conclude that the semantic information of different modalities has a significant impact on the BERT score. Specifically, the “Desc. Only” scheme has the least semantic information, and DJSCC-RGB is heavily affected by low SNR. “Sketch+Desc.” and “Sketches+Desc.” schemes maintain basic control over structural and action semantics, so the BERT score is relatively high. The “First Frame+Desc.” scheme conveys richer semantic information and consequently achieves the highest BERT score.
@@
Fig. 5 presents the PSNR performance of different schemes for various SNRs. Compared with other performance metrics, PSNR focuses primarily on pixel-level reconstruction quality. As depicted in Fig. 5, methods lacking strong semantic control capabilities usually struggle to achieve pixel level consistency, resulting in the lower PSNR scores. Although DJSCC-RGB may appear blurry, it achieves higher pixel similarity than “Sketch+Desc.” and “Sketches+Desc.” schemes. However, schemes using the first frame as semantic information still perform effectively on the PSNR, demonstrating the robust pixel-level accuracy.
@@
As illustrated in Fig. 6, SSIM assesses the structural differences in video content. The “Desc. Only” method lacks structured semantic controls and has the lowest SSIM score as shown in Fig. 6, indicating its poor performance in structural integrity. In contrast, methods utilizing sketches exhibit a gradual increase in SSIM scores with improving SNR, reflecting the enhanced ability to convey structural semantic information. This upward trend in SSIM scores highlights the capability of “Sketch+Desc.” and “Sketches+Desc.” schemes to improve structural fidelity in video reconstruction, thus enhancing the overall quality and viewing experience. Moreover, the “First Frame+Desc.” scheme, which utilizes the most comprehensive semantic information, exhibits the strongest performance in terms of structural semantics. This is reflected in its consistently high SSIM scores in all SNR levels, highlighting its effectiveness in preserving both the integrity and detail of video structure, thereby providing a compelling case for its use in scenarios that demand high fidelity in semantic and structural aspects.

### v2

@@ Abstract
Despite significant advancements in traditional syntactic communications based on Shannon’s theory, these methods struggle to meet the requirements of 6G immersive communications, especially under challenging transmission conditions. With the development of generative artificial intelligence (GenAI), progress has been made in reconstructing videos using high-level semantic information. In this paper, we propose a scalable generative video semantic communication framework that extracts and transmits semantic information to achieve high-quality video reconstruction. Specifically, at the transmitter, description and other condition signals (e.g., first frame, sketches, etc.) are extracted from the source video, functioning as text and structural semantics, respectively. At the receiver, the diffusion-based GenAI large models are utilized to fuse the semantics of the multiple modalities for reconstructing the video. Simulation results demonstrate that, at an ultra-low channel bandwidth ratio (CBR), our scheme effectively captures semantic information to reconstruct videos aligned with human perception under different signal-to-noise ratios. Notably, the proposed First Frame+Desc. scheme consistently achieves CLIP score exceeding 0.92 at CBR = 0.0031 for SNR >> 0 dB. This demonstrates its robust performance even under low SNR conditions.
@@ I Introduction
On the other hand, advancements in generative artificial intelligence (GenAI) not only enable the creation of realistic, high-quality content but also facilitate its efficient transmission through semantic communication innovations, thereby optimizing bandwidth usage without compromising on quality. Early GenAI focused on probabilistic graphical models, such as hidden Markov models [7], which generated data with limited quality and diversity. Variational autoencoders (VAEs) marked a breakthrough by producing more complex and realistic data [8, 9], but they faced distribution mismatch issues. To address this, generative adversarial networks (GANs) were proposed [10, 11], optimizing the generator and discriminator through adversarial loss to improve data quality. GANs have been used after DJSCC at the receiver to enhance perception quality [12]. However, GANs are prone to instability, such as mode collapse, resulting in a lack of diversity in the generated samples.
Recently, many large models using diffusion techniques have appeared, and they can generate data based on conditional information. [13, 14]. For example, the prompt “Panda plays ukelele at home” can be used to generate a short video whose semantic content aligns with the description provided in the prompt [15]. Due to the generative models’ advantages in accurate grasp of semantic conditions and the substantial improvement of content authenticity, the authors in [16] and [17] focused on extracting and transmitting compressed semantic features, such as text and edge maps, to guide the diffusion process at the receiver. They demonstrated that even at ultra-low compression rates (less than 0.1 bit per pixel), essential semantics can still be preserved at the receiver. However, these works focus on image transmission, and the research on video transmission is still lacking. The authors in [18] and [19] explored the use of diffusion models to reconstruct inter-frame information in video frames, focusing primarily on video compression in error-free channels.
@@
Therefore, a significant research gap exists in effectively utilizing diffusion models for video reconstruction under practical wireless transmission channels. Delivering video semantics as conditional signals for diffusion models to achieve high semantic scores remains an open challenge.
To fill these gaps, we design a generative video semantic communication (GVSC) framework for transmitting video semantics and reconstructing videos, whereby GenAI large models with the diffusion techniques are utilized to generate video aligned with human perception from video semantics of sketch sequences and text description. Our contributions are summarized as follows:
@@
1) We propose a novel GVSC framework for video reconstruction at ultra-low channel bandwidth ratios (CBR). At the transmitter, our semantic extractor identifies key visual and textual semantics from the video. The visual information is transmitted via DJSCC, while the textual information uses turbo coding with quadrature amplitude modulation (QAM). These schemes include critical details such as object location, color, size, and actions. Unlike traditional schemes, GVSC leverages a pre-trained GenAI model to reconstruct video content from condensed semantic data. At the receiver, a GenAI diffusion model integrates this information for effective video reconstruction. The robustness of GVSC ensures effectiveness even in resource-constrained environments, performing well at CBR as low as $10^{-2}$.
@@ II-B2 Joint Source Channel Coding for Visual Modality
$\mathcal{L}(x,x_{0})=k\cdot\text{MSE}(x,x_{0})+(1-k)\cdot\text{LPIPS}(x,x_{0}),$ (1)
@@ II-C2 Diffusion Process
After obtaining the semantic embedding, the video generator utilizes these information as conditions to reconstruct the video at the receiver. During the generation process, the diffusion model $\epsilon_{\theta}$ combines the embeddings of the visual semantics and the video description semantics to perform denoising. Assuming $z_{t}$ is the latent representation at time step $t$, the generation process is as follows:
@@ III-A Transmission Strategy and Model Configuration
We calculate the widely-adopted channel bandwidth ratio as $\text{CBR}=\frac{K}{C_{\text{in}}HWF}\quad\text{symbols/pixel}$ [23, 6]. Here, $K$ denotes the total number of symbols transmitted over the wireless channels, while $C_{\text{in}}$, $H$, $W$, and $F$ correspond to the number of input channels, height, width, and video frames, respectively.
@@
Sketch+Desc.: To enhance the clarity of visual semantics during transmission, we opt for transmitting only the initial frame’s sketch, thereby concentrating resources on its quality. An initial sketch of the video is transmitted with video description. The sketch is encoded using the DJSCC optimized for sketch transmission. Stable Diffusion 3.5 [26] is used to reconstruct the first frame of the video, after which Open-Sora [27] reconstruct the whole video. We configure the DJSCC-sketch and train the model over an SNR range from 0 to 10 dB. Testing is performed across all the SNRs within this range. The total number of symbols required for this strategy is given by $K=1.5N_{d}+N_{s}$ where $N_{s}$ represents the number of transmitted symbols for the sketch, and $1.5N_{d}=1147.56$ symbols are required for the video description. The output size of the DJSCC encoder is $32\times 32\times 2$. Therefore, the resource blocks allocated for the sketch data are calculated as $\frac{32\times 32\times 2}{2}=1024$ i.e., resource blocks for sketch. Thus, the total number of resource blocks required for the Sketch+Desc. scheme is:$K=1147.56+1024=2171.56$. Consequently, the CBR for this scheme is $0.001$.
First Frame+Desc.: For acquiring more detailed semantic information, we choose to transmit the first frame with video descriptions, which are transmitted using DJSCC. Open-Sora then utilizes these rich semantic inputs to reconstruct the video. This scheme combines a high-quality visual anchor with textual semantics to detail dynamic content, ensuring the high-quality video reconstruction. For this scheme, we have specifically trained the DJSCC to optimize the transmission of the first frame. The video description is transmitted as Desc. Only scheme. This scheme achieve a average CBR of $0.0031$.
@@ III-B Evaluation Metrics
The GVSC Framework adopts the following metrics to evaluate semantic similarity. CLIP score: The CLIP score [28], widely adopted due to its training on large image-text datasets [29, 30, 31, 32], effectively captures high-level semantic features and is suitable for assessing semantic similarity. We use the average CLIP score to compare each generated frame with its corresponding original frame. It effectively measures the semantic similarity of each video frame. The CLIP score focuses on frame-level semantic similarity, ensuring each generated frame matches its original counterpart. BERT score: We design a video semantic consistency check scheme. To assess overall semantic similarity, we use Video-LLaVA to generate captions for both videos and evaluate their similarities with a pretrained BERT model [33]. Given Video-LLaVA’s generative diversity, we sample captions 3 times for each video and average the scores for the overall semantic score. The BERT score evaluates the video-level semantic consistency by comparing generated captions, measuring how well the generated video captures the broader narrative and context.
PSNR and SSIM [34]: The peak signal-to-noise ratio (PSNR) and structure similarity index measure (SSIM) are traditional video quality metrics, providing insights into pixel and structure-level fidelity. They measure the peak SNR and structural similarity, respectively.
@@ IV-A1 Dataset
We use the WebVid dataset [35] for both training and testing. WebVid contains Internet-crawled videos with captions covering diverse scenes such as human actions and natural landscapes. We use the captions as semantic descriptions, with each video containing $F=8$ frames at a resolution of $256\times 256$ ($H=W=256$). For training the transmission model, we extract over 11,000 sketches using PiDiNet [36], and 129 videos for testing.
@@ IV-A2 Simulation Parameters
For transmitting sketches, we train the DJSCC for 200 epochs using the Adam optimizer and batchsize is 16. The learning rate is $1\times 10^{-4}$. For transmitting first frame, we adjust the learning rate to $1\times 10^{-5}$. Following the human assessment simulations, we set $k=0.3$ in (1). We train our transmission model over a SNR range of 0 to 10 dB and test all models within this range. Due to the excessive number of parameters, we freeze the GenAI model. We set the frame rate to 4, the inference time-step to 50, and the random seed to 7777.
@@ IV-B Comparison Schemes
H.264/H.265+LDPC: We use H.264 and H.265 for source coding, with LDPC codes and QAM modulation for channel coding. LDPC rates and modulation orders are adapted to the SNR, while keeping CBR = 0.005 to ensure fair comparison. Detailed configurations are listed in Table I.
TABLE I: LDPC and modulation configurations under different SNRs
CBR 0 dB 2 dB 4 dB
0.005 1/3 LDPC+4QAM 1/2 LDPC+4QAM 2/3 LDPC+4QAM
CBR 6 dB 8 dB 10 dB
0.005 3/4 LDPC+4QAM 1/2 LDPC+16QAM 2/3 LDPC+16QAM
DJSCC-RGB: We use DJSCC with three layers of upsampling and downsampling modules. We maintain the original channel setting (3 channels) to transmit the RGB frames of the videos. The average CBR of this scheme is $0.005$.
@@ IV-C Simulation Results
Visual comparisons of different transmission schemes at SNR = 10 dB and 0 dB are shown in Fig. 2. The first column on the left displays the original video frames (the 1st and the 8th frames). Subsequent columns present visual results of various methods at SNR = 10 dB. The right part highlights the DVST and First Frame+Desc. schemes at SNR = 0 dB, both maintaining a CBR of $10^{-2}$. H.264+LDPC and H.265+LDPC are able to retain structural details in high SNR. DVST shows the effective transmission at SNR = 10 dB with high visual quality, while DJSCC-RGB undergoes significant visual degradation due to ultra-low CBR, showing highly distorted frames. The Desc. Only scheme preserves basic semantic information but loses structural details, leading to relatively poor video reconstruction. Conversely, the Sketch+Desc. and Sketches+Desc. schemes excel in retaining both structural and textual semantics at lower CBR, as reflected in descriptions such as “A boy is eating corn”. The First Frame+Desc. scheme allocates more resources to the first frame, ensuring semantic consistency and excelling in visual clarity even at SNR = 0 dB, as highlighted in the red circles.
@@
Figure 7: CLIP score for various SNRs with different k in Sketch+Desc. scheme.
In Figs. 445 and 6, the red circles in these figures correspond to different LDPC code rates and modulation schemes used in our experiments. Fig. 4 shows the CLIP scores of different schemes across SNR levels. With adaptive coding, H.264+LDPC fails to work at 0 dB, and H.265+LDPC still underperforms compared to First Frame+Desc. under low SNR conditions. DJSCC-RGB and Desc. Only are constrained by the lack of visual context, while Sketch+Desc. and Sketches+Desc. benefit from structural information and show improvements. DVST improves gradually with increasing SNR, but is eventually surpassed by H.265+LDPC at higher SNRs. Notably, First Frame+Desc. consistently achieves the highest performance in all SNRs, demonstrating the robustness of GVSC framework under challenging channel conditions.
@@
Fig. 4 presents BERT score performance of different schemes for various SNR conditions. Similar to CLIP score, H.264+LDPC has difficulty in reliably conveying the at SNR == 0 dB. The BERT scores for Sketch+Desc. and Sketches+Desc. schemes are lower than H.264+LDPC and H.265+LDPC because traditional schemes transmit detailed visual information closely related to the original video, leading to higher text similarity scores. When generating descriptions, the visual model might include non-critical details such as hair color or eye direction, which are not always relevant in the video generation context. Since BERT focuses on specific textual details, even slight discrepancies can result in lower scores. We can conclude that the semantic information of different modalities has a significant impact on the BERT score. Specifically, the Desc. Only scheme has the least semantic information, and DJSCC-RGB is heavily affected by low SNR. Sketch+Desc. and Sketches+Desc. schemes maintain basic control over structural and action semantics, so the BERT score is relatively high. When SNR exceeds 2 dB, H.264+LDPC, H.265+LDPC, and First Frame+Desc. achieve similar BERT scores, indicating that these scheme provide comparable support for coarse-grained caption generation under moderate-to-good channel conditions.
@@
Fig. 5 presents the PSNR performance of different schemes for various SNRs. Compared with other performance metrics, PSNR focuses primarily on pixel-level reconstruction quality. As depicted in Fig. 5, methods lacking strong semantic control capabilities usually struggle to achieve pixel level consistency, resulting in the lower PSNR scores. Although DJSCC-RGB may appear blurry, it achieves higher pixel similarity than Sketch+Desc. and Sketches+Desc. schemes. However, schemes using the first frame as semantic information still perform effectively on the PSNR, demonstrating the robust pixel-level accuracy.
@@
As illustrated in Fig. 6, SSIM assesses the structural differences in video content. The Desc. Only method lacks structured semantic controls and has the lowest SSIM score as shown in Fig. 6, indicating its poor performance in structural integrity. In contrast, methods utilizing sketches exhibit a gradual increase in SSIM scores with improving SNR, reflecting the enhanced ability to convey structural semantic information. This upward trend in SSIM scores highlights the capability of Sketch+Desc. and Sketches+Desc. schemes to improve structural fidelity in video reconstruction, thus enhancing the overall quality and viewing experience. Moreover, the First Frame+Desc. scheme, which utilizes the most comprehensive semantic information, exhibits the strongest performance in terms of structural semantics. This is reflected in its consistently high SSIM scores in all SNR levels, highlighting its effectiveness in preserving both the integrity and detail of video structure, thereby providing a compelling case for its use in scenarios that demand high fidelity in semantic and structural aspects.